Solve the system by addition. If a unique solution does not exist, state whether the system is inconsistent or dependent.
x+5y=10
-2x-10y= -20
Show your solution below:
What is the solution for this system?
What type of system is this?











Found 2 solutions by gems, stanbon:
Answer by gems(12)

from x+5y=10, use substitution method x= 10-5y
using 2nd equation -2x+10y=-20,
-2(10-5y)+10y=-20
-20+10y+10y=-20
20y=-20+20
20y=0
y=0
ify=0,
x+5y=10
x+5(0)=10
x+0=10
x=10
So the solution is x= 10, y=0
the equation is consistent

Answer by stanbon(75887)

Solve the system by addition. If a unique solution does not exist, state whether the system is inconsistent or dependent.
x+5y=10
-2x-10y= -20
----------------
Multiply thru the 1st equation by -2 to get:
-2x-10y = -20
---------
Subtract that from the 2nd equation to get:
0=0
-----------
That means one of the equations is a multiple of the other.
That means the system is "dependent".
Cheers,
Stan H.
Show your solution below:
What is the solution for this system?
the solution is the set of points on the line x + 5y = 10
---------------
What type of system is this? dependent
-------------